# Can In-context Learning Really Generalize to Out-of-distribution Tasks?

### v1

@@ 6 Conclusion
In this work, we empirically find that Transformers struggle to generalize beyond the pretraining function classes when given downstream in-context examples of OOD tasks. Instead, ICL tries to seek a near-optimal solution within the pretraining function classes. However, ICL performs well in retrieval tasks where the shift in the input-label mapping is only caused by replacing the in-context label tokens with new ones while the underlying function distribution retains. We also examine ICL’s performance on OOD tasks after pretraining on multiple tasks. Our theoretical and empirical analysis reveals ICL’s preference for low-test-error functions, i.e., ICL tends to implement pretraining function classes with low test error in the test context. This finding, alongside previous work (Lin & Lee, 2024), highlights two key factors that determine how ICL will implement the prediction function based on the testing context and pretraining tasks: the distance between the training and testing input distributions, and the ability of a pretraining function to solve the test task.
7 Limitations
1) Most experimental results are based on a GPT-2 model pretrained on a limited set of mathematical functions. It is challenging to assess whether modern large-scale language models like GPT-4 and Claude 3 Opus face similar difficulties in generalizing beyond their pretraining corpus, given the vast range of tasks and content in their pretraining data. Nevertheless, our findings highlight the limitations of ICL for smaller models like Llama2-7b. 2) The models are trained on ICL data, while real-world LLMs are trained autoregressively. However, the ICL pretraining objective is also next-token prediction, so there is no clear gap between these two pretraining objectives.
8 Reproducibility
In the main text and Appendix A, we’ve stated all setups for reproducing our experimental results. For the theoretical part, we’ve included the assumptions (Assumption C.2) and proofs in Appendix C.
Appendix A Experimental Details
A.1 Experimental Details in Section 2 and Section 4.2
@@ Appendix A Experimental Details
A.2 Experimental Details in Section 3.1
@@
A.3 Experimental Details of Section 3.3
@@
We uniformly sample 1000 word vectors $\bm{x_{i}}\in\mathbb{R}^{d}$ from the word embedding $E\in\mathbb{R}^{N\times d}$ of the pretrained Llama-2-7b, where $N=32000$ and $d=4096$. Then we generate a task weight $\bm{w}\in\mathbb{R}^{d^{\prime}\times C}$ that only takes the first $d^{\prime}$ dimensions of $\bm{x}_{i}$ (denoted as $\bm{x}_{i}[:d^{\prime}]$) to compute a probability distribution over $C$ classes: $p_{i}=\bm{x}_{i}[:d^{\prime}]^{\top}\bm{w}\in\mathbb{R}^{C}$. Thus, each $\bm{x}_{i}$ is mapped to one of the $C=10$ classes. Next, we sample the label vectors $y_{i}\in\mathbb{R}^{d}$ from $E$ according to the index of maximum probability: $y_{i}=E_{\arg\max_{j}p_{i}[j]+s}$, where $s\in\mathcal{N}+$ is a offset. We set $d^{\prime}=30\ll d=4096$ to reduce the complexity of the task. Hence, $\bm{x}_{i}$ are classified into $C$ labels words $E[s:s+C]$. The predicted token of $\bm{x}_{i}$ is computed as: $\arg\max_{j}\hat{p}_{i}[j],~{}j\in\{s,s+1,...,s+C-1\}$, where $\hat{p}_{i}$ is the output of the last linear layer of Llama-2-7b given $\bm{x}_{i}$.
Appendix B Additional Theoretical Details
B.1 Additional Details of Lemma 2.1
In this section, we provide the necessary details of the theoretical setting of Zhang et al. (2023).
@@ Appendix B Additional Theoretical Details
The linear self-attention (LSA) model considered in the Theorem 4.2 of Zhang et al. (2023) (Lemma 2.1) is defined as follows:
@@
Now we present the assumption on the attention weights of the linear-attention model in Lemma 2.1.
Assumption B.1.
(Assumption 3.3 in Zhang et al. (2023), initialization). Let $\sigma>0$ be a parameter, and let $\Theta\in\mathbb{R}^{d\times d}$ be any matrix satisfying $\left\|\Theta\Theta^{\top}\right\|_{F}=1$ and $\Theta\Lambda\neq 0_{d\times d}$. We assume
@@
Appendix C Proof of Theorem 4.4
We restate Theorem 4.4 as the Theorem C.1 below, and present the assumption it depends on.
Theorem C.1.
(ICL prediction favors the pretraining function with low error on the context) Given the context $S_{k}$, if the empirical risk of implementing a function of the pretraining task $\alpha$ is less than that of $\beta$, i.e., $\frac{1}{T}\sum_{t=1}^{T}|\bm{w}_{\beta}\bm{x}_{i}-y_{i}|^{2}-|\bm{w}_{\alpha}\bm{x}_{i}-y_{i}|^{2}\geq 0$, then, under some mild Assumptions C.2, we have $\Psi_{\bm{w}}(\alpha,\beta)\geq 0$.
@@
In the Appendix H.1 of Lin & Lee (2024), they have proved that when the context length $T\rightarrow\infty$, under the first condition in Assumption C.2, $\lim_{T\rightarrow\infty}\Psi_{\mu}(\alpha,\beta)=\geq 0$.
@@
where equation $(a)$ is due to the third condition in Assumption C.2, equation $(b)$ is by applying the Sherman–Morrison formula. Since $\frac{\delta_{w}}{1+\delta_{w}\sum_{i=1}^{T}}\geq 0$, to prove that $\Psi_{\bm{w}}(\alpha,\beta)\geq 0$, we only need to show that
@@
Further, we can derive that the term $\frac{1}{T}\sum_{i=1}^{T}\|\bm{w}_{\beta}-\bm{x}_{i}y_{i}\|^{2}_{\bm{x}_{i}\bm{x}_{i}^{T}}-\|\bm{w}_{\alpha}-\bm{x}_{i}y_{i}\|^{2}_{\bm{x}_{i}\bm{x}_{i}^{T}}$ below is non-negative by using the condition 2 in Assumption C.2:
@@
where the inequality $(c)$ holds since according to the condition 2 in Assumption C.2, $\bm{x}_{i}\bm{x}_{i}^{T}-\bm{I}$ is positive semi-definite, and the inequality $(d)$ holds since the population downstream risk of $\alpha$ is lower than $\beta$. Therefore, to prove inequality (8), we just need to prove that the l.h.s. of inequality (8) multiplying $\frac{1}{T}$ is not less than $\frac{1}{T}\sum_{i=1}^{T}\|\bm{w}_{\beta}-\bm{x}_{i}y_{i}\|^{2}_{\bm{x}_{i}\bm{x}_{i}^{T}}$ in Equation (9):
@@
applying the condition 4 in Assumption C.2, we get the final conclusion.

### v3

@@ 6 Conclusion
In this work, we empirically find that Transformers struggle to generalize beyond the pretraining function classes when given downstream in-context examples of OOD tasks. Instead, ICL tries to seek a near-optimal solution within the pretraining function classes. We further investigate the widely observed capability of ICL to perform classification. We reveal that it is a composition of ID prediction and retrieval rather than an OOD generalization ability. We also examine ICL’s performance on OOD tasks after pretraining on multiple tasks. Our theoretical and empirical analysis reveals ICL’s preference for low-test-error functions, i.e., ICL tends to implement pretraining function classes with low test error in the test context. This finding, alongside previous work (Lin & Lee, 2024), highlights two key factors that determine how ICL will implement the prediction function based on the testing context and pretraining tasks: the distance between the training and testing input distributions, and the ability of a pretraining function to solve the test task.
Appendix A Comparison with Related Works and Additional Discussions
A.1 The Capability of ICL to Learn New Tasks
Besides studies indicating that ICL can learn new weights of linear regression (Garg et al., 2022; Raventós et al., 2023; Zhang et al., 2023a; Akyürek et al., 2023), other research has found that LLMs can tackle tasks that are unlikely to have been encountered during pretraining. For example, Pan (2023) showed that LLMs perform better than random guessing on classification tasks with meaningless labels. Kossen et al. (2024) demonstrate that ICL can identify authorship based on writing style in private communication messages not included in the pretraining corpus. Additionally, Vacareanu et al. (2024) found that large-scale LLMs can learn various linear and non-linear functions from context. We argue that these findings do not contradict our work. While the LLMs may not have seen exactly the same tasks, there is no guarantee that they haven’t encountered tasks from a similar distribution in their pretraining corpus. For instance, the LLMs could have been pretrained on a corpus containing authorship identification tasks or on statistical data encompassing different functions. Our work does not claim that ICL cannot generalize to new task instances; rather, it highlights the limitation in generalizing to an unseen input-label distribution. Additionally, Yadlowsky et al. (2023) finds that ICL struggles to generalize to testing function classes that are unseen during training (e.g., convex combinations or extreme versions of the pretraining functions). They didn’t delve into how ICL behaves on OOD data, while we reveal that it implements the pretraining functions.
A.2 The Algorithm-selection Mechanism of ICL
Recent works by Bai et al. (2023); Wang et al. (2024) have uncovered the algorithm selection phenomenon, demonstrating that Transformers pretrained on both linear regression and classification tasks perform well when presented with the context of either task during ICL inference. Theoretically, they show that a Transformer with specific parameters can achieve algorithm selection. Yadlowsky et al. (2023) empirically found that ICL selects the optimal pretraining function class after observing in-context examples from a function class present in the pretraining data mixture. However, the algorithm selection experiments in these studies are limited to scenarios where the test functions are among the training functions. In this work, we empirically and theoretically demonstrate that the algorithm selection phenomenon broadly occurs when given downstream context from arbitrary function classes. To the best of our knowledge, we are the first to reveal the factors that determine the selection process.
A.3 The Bayesian-optimal Perspective for Understanding ICL
Many studies have found that ICL makes Bayes-optimal predictions (Xie et al., 2022; Wies et al., 2024; Zhang et al., 2023b; Lin & Lee, 2024). However, these works have certain limitations that may reduce their practical applicability in predicting ICL behavior in general scenarios. 1) Limited empirical verification. Wies et al. (2024) and Zhang et al. (2023b) lack empirical verification of their theory on real deep Transformer models; 2) Limited theoretical settings: in-distribution tasks. Wies et al. (2024) assumes the downstream tasks are components of the pretraining distribution; Xie et al. (2022) assumes that the latent concept of the test task $\theta^{*}$ is within the pretraining task set $\Theta$; In Lin & Lee (2024), the training and testing tasks are all linear regression with weights sampled from Gaussian distribution. 3) Limited implications of the theoretical results: although Xie et al. (2022); Zhang et al. (2023b) prove that ICL can infer a task concept $\theta$ based on the downstream test context $S_{test}$, they don’t reveal how $S_{test}$ concretely affects the posterior distribution $P(\theta|S_{test})$ of the latent task concept $\theta$ inferred by the model that determines the downstream ICL prediction, especially when the true downstream task $\theta^{*}$ is OOD. Our work verifies and extends previous findings to a more general setting by using real deep Transformers and evaluating ICL on OOD tasks that significantly differ from the training tasks. For the first time, we also reveal how the interaction between the downstream distribution and the pretraining distribution affects ICL predictions (see Section 5).
In contrast, Raventós et al. (2023) claim that ICL can exhibit non-Bayesian properties. They empirically demonstrate that when given sufficiently diverse pretraining tasks (linear regression vectors), ICL can outperform the Bayesian estimator on a new test distribution. However, the distributional shift in their setup might not be substantial enough to show that ICL can truly adapt to a new downstream distribution, which is considered to be “non-Bayesian” by Raventós et al. (2023). In their setting, both the test and training vectors are sampled from the standard Gaussian distribution, and the only source of ”distributional shift” comes from the finite size of the training set, which can only partially reflect the test distribution. Our work refutes their findings by showing that when the test distribution is significantly shifted, increasing the number of ID tasks may not help ICL generalize to it.
A.4 The Bayesian Interpretation for Our Empirical Findings
Although current Bayesian theories for ICL are too vacuous to predict the performance of deep Transformers on real OOD tasks (see A.3), the Bayesian framework shows promise as a potential lens for interpreting our empirical findings. Here we provide some intuitive interpretations for the findings in Section 3, 4, and 5 from a Bayesian perspective.
Consider the predicted distribution $p_{\theta}(\bm{y}_{T}|\bm{x}_{1:T})$ given by a pretrained model $\theta$. If we assume that ICL makes Bayesian-like predictions over the test context as (Xie et al., 2022; Wies et al., 2024; Zhang et al., 2023b; Lin & Lee, 2024) suggested, then the model will first infer a task concept $\phi$ based on the given context $\bm{x}_{1:T-1}$ and predict $\bm{y}_{T}$ using $\phi$ and $\bm{x}_{1:T}$, i.e.,
$p_{\theta}(\bm{y}|\bm{x}_{1:T})=p_{\theta}(\bm{y}|\bm{x}_{1:T},\phi)p_{\theta}(\phi|\bm{x}_{1:T-1})$ (3)
To explain the results in Section 3 and Section 5: since the true downstream task $\phi^{*}$ is unseen during pretraining, the inferred posterior distribution $p_{\theta}(\phi|\bm{x}_{1:T-1})$ assigns probability mass only to tasks $\phi$ within the pretraining distribution that maximize $p_{\theta}(\bm{y}|\bm{x}_{1:T})$. This accounts for why ICL can only make in-distribution predictions, as shown in Section 3, and why ICL prefers pretraining priors with low test error and input distributions similar to those in the test context. Once a task $\phi$ seen during pretraining is identified as best fitting the test context $\bm{x}_{1:T-1}$, the model refines its predictions based on this context. This refinement corresponds to the factor $p_{\theta}(\bm{y}|\bm{x}_{1:T},\phi)$, explaining how ICL optimizes predictions within its pretraining distribution.
In Section 4, the underlying task concept $\phi$ acts as a similarity metric that allows the model to retrieve examples from the context that align with the query. Training on more abstract labels improves the model’s ability to estimate a more accurate $\phi$. When the test task is ID, even with OOD labels, ICL can succeed by leveraging the learned $\phi$ to predict the true label. It accomplishes this by retrieving an example $\bm{x}_{i}$ from the context that is similar to the query under the $\phi$ metric. However, when the underlying task $\phi^{*}$ is OOD, the model fails because the learned similarity metric no longer applies effectively.
A.5 Discussion of the Setup of Our Theory
Notably, our theoretical result in Section 5.1 doesn’t assume a Transformer architecture, while previous theoretical works of understanding ICL often adopt Transformers with oversimplified assumptions on their parameters or structures (Ahn et al., 2023; Zhang et al., 2023a; Huang et al., 2023; Collins et al., 2024). Additionally, our analysis shows that models pretrained on the ICL tasks can implement algorithm selection during ICL inference following Lin & Lee (2024). In contrast, prior work on algorithm selection (Bai et al., 2023) only shows that a specific set of parameters in a simplified ReLU Transformer can enable algorithm selection. However, the parameter construction is complex and somewhat tricky, and there is no theoretical or experimental guarantee that Transformers exhibiting algorithm selection will necessarily implement these parameters.
Appendix B Additional Experimental Results
B.1 Understanding the Effect of Training on More Diverse Retrieval Tasks from the Attention Scores
(a) PT $s\sim\mathcal{U}(50,150)$
(b) PT $s\sim\mathcal{U}(50,250)$
(c) PT $s\sim\mathcal{U}(50,450)$
Figure 10: The matching score of all attention heads of models trained on the retrieval task. “PT” denotes “pretrained on”. Each subfigure corresponds to a different pretrained model. The model of (c) exhibits more heads with high matching scores, which is also the most performant model in the retrieval task in Figure 3.
B.2 The Synthetic Word Classification is Not That Hard to Solve If It’s In Distribution
To show the failure in the synthetic word classification in Section 4.3 is mainly due to its OOD nature rather than it’s intrinsically too hard to learn, we train a GPT-2 to perform the same task as in Section 4.3. In this task, the $\bm{x}_{i}$ and $\bm{y}_{i}$ are generated in the same way as Section 4.3. The only modification is that we use a smaller predefined vector embedding $E^{\prime}\in\mathbb{R}^{10000\times 20}$ ($E_{llama}\in\mathbb{R}^{32000\times 4096}$ in the experiment in Section 4.3). The results in Figure 11 show that when $\bm{W}$ has been encountered during pretraining, ICL can well address this task.
Figure 11: Test error of the GPT-2 trained and evaluated on the same synthetic OOD word classification task as in Section 4.3.
B.3 Evaluating the Synthetic OOD Classification Task on Llama-2-7B
We also evaluate Llama-2-7B on the same OOD word classification task and the retrieval task as in Section 4.3. Figure 12 shows the same observations as in Figure 6 that the LLM can well address the retrieval task but fails to learn the OOD function $\bm{W}$. In this experiment, we set the length of the repeating sequence to be 10. We can observe that the accuracy of retrieval rapidly increases after seeing 10 in-context examples. This demonstrates that learning novel functions from the context is challenging for real-world pretrained LLMs, but the LLMs are good at retrieving.
B.4 Will Generalization Capabilities Emerge from Increasing the Number of Training Tasks?
Recent work by Raventós et al. (2023) empirically demonstrates that when both the training and test tasks are linear regression, and the number of training vectors exceeds a certain ”task diversity threshold” (approximately $2^{14}\sim 2^{15}$), ICL can generalize from a finite training set sampled biasedly from $\mathcal{N}(0,1)$ to the test distribution $P_{\text{test}}=\mathcal{N}(0,1)$ (see Appendix A.3 for details). We investigate whether similar phenomena persist for test tasks with larger distributional shifts. We train models using varying numbers of linear regression vectors and evaluate them on quadratic and ReLU neural network regression tasks. In Figure 13, we find that training on a vast number of in-distribution (ID) functions does not yield any improvements, providing further evidence that ICL may struggle to achieve OOD generalization.
(a) quadratic regression
(b) ReLU NN regression
Figure 13: The ICL test error of models trained on different numbers of linear regression vectors. Even if the number of training vectors (up to $1,000,000\approx 2^{20}$) surpasses the threshold ($2^{14}\sim 2^{15}$) reported by Raventós et al. (2023), no model exhibits generalization to OOD function classes.
Appendix C Experimental Details
C.1 Experimental Details in Section 3.1 and Section 5.2
@@ Appendix C Experimental Details
C.2 Experimental Details for Section 3.2
For the reversed-label experiment, we choose four tasks: Antonym, Capital-country, English-French, and English-German. The original datasets are adopted from Todd et al. (2024). The top-1 accuracy is computed as follows: compute the top-1 accuracy for each token predicted by the model, based on the token length of the ground-truth label word. For each context length, we compute the average accuracy over 128 test examples.
C.3 Experimental Details for Section 4.1
@@
C.4 Experimental Details for Section 4.3
@@
We uniformly sample 1000 word vectors $\bm{x}_{i}\in\mathbb{R}^{d}$ from the word embedding $E\in\mathbb{R}^{N\times d}$ of the pretrained Llama-3-8B, where $N=128256$ and $d=4096$. Then we sample a task weight $\bm{W}\in\mathbb{R}^{d^{\prime}\times C}$ from standard Gaussian distribution that only takes the first $d^{\prime}$ dimensions of $\bm{x}_{i}$ (denoted as $\bm{x}_{i}[:d^{\prime}]$) to compute a probability distribution over $C$ classes: $p_{i}=\bm{x}_{i}[:d^{\prime}]^{\top}\bm{W}\in\mathbb{R}^{C}$. Next, we set the label vectors $\bm{y}_{i}=E_{\arg\max_{j}p_{i}[j]+s}\in\mathbb{R}^{d}$, where $s=10000$ is a offset. We set $d^{\prime}=30\ll d=4096$ to reduce the complexity of the task. Hence, $\bm{x}_{i}$ are classified into $C$ labels words $E[s:s+C]$. The predicted token of $\bm{x}_{i}$ is computed as: $\arg\max_{j}\hat{p}_{i}[j],~{}j\in\{s,s+1,...,s+C-1\}$, where $\hat{p}_{i}$ is the output of the last linear layer of Llama-3-8B given $\bm{x}_{i}$.
C.5 Experimental Details for Section 5.3
In this section, we present some details about the setups for the ambiguous classification task. The label mapping rule is presented in Table 2. For each context length, we compute the average accuracy over 128 test examples.
Table 2: Experiment setting of Figure 8(a) and Figure 8(b). “PT” and “DS” are short for “pretraining” and “downstream”, respectively.
Classification criterion Original labels Labels presented in the context
sentiment “positive” “RqF”
“neutral” “IwZ”
“negative” “SdK”
type “science” “RqF”
“sports” “IwZ”
“arts” “SdK”
location “Asia” “RqF”
“Europe” “IwZ”
“Africa” “SdK”
Prompt examples. Here we present some in-context examples of the input prompt of using different classification criteria.
Using “sentiment” as the classification criterion. Q: The groundbreaking discovery made by Japanese scientists has revolutionized renewable energy. A: RqF gray# Original label: “positive” Q: A chess championship occurred in Russia, featuring players from around the continent. A: IwZ gray# Original label: “neutral”
Using “type” as the classification criterion. Q: A regional basketball league was formed in Kenya to promote the sport locally. A: IwZ gray# Original label: “sports” Q: The breathtaking architectural exhibition in Dubai left visitors absolutely awestruck. A: SdK gray# Original label: “arts”
Using “location” as the classification criterion. Q: A scientific paper from Finland explores new methodologies in data analysis. A: IwZ gray# Original label: “Europe” Q: An astronomy workshop was conducted in Ethiopia for students interested in space. A: SdK gray# Original label: “Africa”
Accuracy computation. For a given label, the method to calculate top-5 accuracy is as follows: compute the top-5 accuracy for each token predicted by the model, based on the token length of the ground-truth label word. For a classification criterion other than the one selected in the current sequence, to verify whether the model’s prediction distribution across all test samples approaches the label distribution under that criterion, we select the permutation among all possible mappings between original labels and meaningless strings that yields the highest model prediction accuracy to compute the accuracy.
Appendix D Existing Theoretical Evidence Supporting that ICL Makes ID Predictions
One recent work (Zhang et al., 2023a) theoretically proved that a one-layer linear self-attention model (LSA, defined in Appendix D) pretrained on a linear regression task will still implement the linear predictor given downstream in-context examples of arbitrary new function classes, under some assumptions on the initialization of the Transformer weight matrices. We restate the Theorem 4.2 of Zhang et al. (2023a) as Lemma D.1 below:
Lemma D.1.
(Theorem 4.2 of Zhang et al. (2023a), informal) Let $\mathcal{D}$ be a distribution over $(\bm{x},y)\in\mathbb{R}^{d}\times\mathbb{R}$, whose marginal distribution on $x$ is $\mathcal{D}_{x}=$ $\mathcal{N}(0,\Lambda)$. Assume the test prompt is of the form $P=\left(\bm{x}_{1},y_{1},\ldots,\bm{x}_{T},y_{T},\bm{x}_{\text{query }}\right)$, where $\left(\bm{x}_{i},y_{i}\right),\left(\bm{x}_{\text{query }},y_{\text{query }}\right)\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}\mathcal{D}$. The prediction risk on the test query $y_{query}$ of an arbitrary task satisfies:
Lemma D.1 serves as a shred of theoretical evidence that ICL can just implement the pretraining function class, while the role of the context examples is to optimize the model within the pretraining hypothesis space.
Below, we provide the necessary details of the theoretical setting of Zhang et al. (2023a).
@@ Appendix D Existing Theoretical Evidence Supporting that ICL Makes ID Predictions
The linear self-attention (LSA) model considered in the Theorem 4.2 of Zhang et al. (2023a) (Lemma D.1) is defined as follows:
@@
Now we present the assumption on the attention weights of the linear-attention model in Lemma D.1.
Assumption D.2.
(Assumption 3.3 in Zhang et al. (2023a), initialization). Let $\sigma>0$ be a parameter, and let $\Theta\in\mathbb{R}^{d\times d}$ be any matrix satisfying $\left\|\Theta\Theta^{\top}\right\|_{F}=1$ and $\Theta\Lambda\neq 0_{d\times d}$. We assume
@@
Appendix E The Lemmas, Assumption, and Proof for Theorem 5.3
In this section, we will first present the lemmas and assumption that Theorem 5.3 depends on and the provide its proof.
E.1 Lemmas for Theorem 5.3
The lemma below states that the closed-form prediction of the model trained on the pretraining data under Assumption 5.1, given the testing context, remains a Gaussian mixture of the reweighted pretraining task weights:
Lemma E.1.
(Corollary 2 of Lin & Lee (2024), closed-form ICL prediction of the pretrained model) Denote the model $M^{*}$ that minimizes the risk on the pretraining data of Assumption 5.1, i.e., $M^{*}\in\arg\min\frac{1}{T}\sum_{i=0}^{T-1}\underset{S_{i}\sim P_{tr}}{\mathbb{E}}\left[\left\|M\left(\mathcal{S}_{i}\oplus\bm{x}_{i+1}\right)-y_{i+1}\right\|^{2}\right]$, then the prediction on any sequence $\mathcal{S}_{i}\oplus\bm{x}_{i+1}$ by $M^{*}$ is as follows: $M^{*}\left(\mathcal{S}_{i}\oplus\bm{x}_{i+1}\right)=\left\langle\bm{x}_{i+1},\sum_{m=1}^{M}\tilde{\pi}_{m}\tilde{\bm{w}}_{m}\right\rangle$. where $\tilde{\pi}_{m}$, and $\tilde{\bm{w}}_{m}$ depending on both the pretraining task and the downstream context example are given in Lemma 1 of Lin & Lee (2024).
E.2 The Assumption for Theorem 5.3
The assumption below impose some mild requirements on the distribution of the downstream context:
Assumption E.2.
E.3 Proof for Theorem 5.3
Now we restate Theorem 5.3 as the Theorem E.3 below
Theorem E.3.
(ICL prediction favors the pretraining function with low error on the context) Given the context $S_{k}$, if the empirical risk of implementing a function of the pretraining task $\alpha$ is less than that of $\beta$, i.e., $\frac{1}{T}\sum_{t=1}^{T}|\bm{w}_{\beta}\bm{x}_{i}-y_{i}|^{2}-|\bm{w}_{\alpha}\bm{x}_{i}-y_{i}|^{2}\geq 0$, then, under some mild Assumptions E.2, we have $\Psi_{\bm{w}}(\alpha,\beta)\geq 0$.
@@
In the Appendix H.1 of Lin & Lee (2024), they have proved that when the context length $T\rightarrow\infty$, under the first condition in Assumption E.2, $\lim_{T\rightarrow\infty}\Psi_{\mu}(\alpha,\beta)=\geq 0$.
@@
where equation $(a)$ is due to the third condition in Assumption E.2, equation $(b)$ is by applying the Sherman–Morrison formula. Since $\frac{\delta_{w}}{1+\delta_{w}\sum_{i=1}^{T}}\geq 0$, to prove that $\Psi_{\bm{w}}(\alpha,\beta)\geq 0$, we only need to show that
@@
Further, we can derive that the term $\frac{1}{T}\sum_{i=1}^{T}\|\bm{w}_{\beta}-\bm{x}_{i}y_{i}\|^{2}_{\bm{x}_{i}\bm{x}_{i}^{T}}-\|\bm{w}_{\alpha}-\bm{x}_{i}y_{i}\|^{2}_{\bm{x}_{i}\bm{x}_{i}^{T}}$ below is non-negative by using the condition 2 in Assumption E.2:
@@
where the inequality $(c)$ holds since according to the condition 2 in Assumption E.2, $\bm{x}_{i}\bm{x}_{i}^{T}-\bm{I}$ is positive semi-definite, and the inequality $(d)$ holds since the population downstream risk of $\alpha$ is lower than $\beta$. Therefore, to prove inequality (9), we just need to prove that the l.h.s. of inequality (9) multiplying $\frac{1}{T}$ is not less than $\frac{1}{T}\sum_{i=1}^{T}\|\bm{w}_{\beta}-\bm{x}_{i}y_{i}\|^{2}_{\bm{x}_{i}\bm{x}_{i}^{T}}$ in Equation (10):
@@
applying the condition 4 in Assumption E.2, we get the final conclusion.
Appendix F Limitations
1) Most experimental results are based on a GPT-2 model pretrained on a limited set of mathematical functions. It is challenging to assess whether modern large-scale language models like GPT-4 and Claude 3 Opus face similar difficulties in generalizing beyond their pretraining corpus, given the vast range of tasks and content in their pretraining data. Nevertheless, our findings highlight the limitations of ICL in solving challenging tasks for smaller models like Llama-2-7B and Llama-3-8B. 2) The models are trained on ICL data, while real-world LLMs are trained autoregressively. However, the ICL pretraining objective is also next-token prediction, so there is no clear gap between these two pretraining objectives.
Appendix G Reproducibility
In the main text and Appendix C, we’ve stated all setups for reproducing our experimental results. For the theoretical part, we’ve included the assumptions (Assumption E.2) and proofs in Appendix E.